### v6

# Mutation-Guided Unit Test Generation with a Large Language Model

Abstract
Unit tests play a vital role in uncovering potential faults in software. While tools like EvoSuite focus on maximizing code coverage, recent advances in large language models (LLMs) have shifted attention toward LLM-based test generation. However, code coverage metrics—such as line and branch coverage—remain overly emphasized in reported research, despite being weak indicators of a test suite’s fault-detection capability. In contrast, mutation score offers a more reliable and stringent measure, as demonstrated in our findings where some test suites achieve 100% coverage but only 4% mutation score. Although a few studies consider mutation score, the effectiveness of LLMs in killing mutants remains underexplored.
In this paper, we propose MutGen, a mutation-guided, LLM-based test generation approach that incorporates mutation feedback directly into the prompt. Evaluated on 204 subjects from two benchmarks, MutGen significantly outperforms both EvoSuite and vanilla prompt-based strategies in terms of mutation score. Furthermore, MutGen introduces an iterative generation mechanism that pushes the limits of LLMs in killing additional mutants. Our study also provide insights into the limitations of LLM-based generation, analyzing the reasons for live and uncovered mutants, and the impact of different mutation operators on generation effectiveness.
I Introduction
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [49, 26, 16, 4]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [56, 33, 6, 53, 57] and search-based methods [25, 63, 62, 70]—to LLM-based approaches for generating unit tests [23, 72, 34].
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [12, 5, 36], high code coverage does not necessarily imply strong fault detection capability. Other research [12, 3, 24] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
Recent efforts have begun exploring mutation-based evaluation for LLM-generated test cases. For example, Foster et al. [24] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases from existing ones and generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [17] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze why live mutants remain undetected, nor does it consider uncovered mutants, both of which are challenges we address in our research.
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly maximizes the mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [68, 74, 58], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect use of assertion functions, including erroneous invocations and inappropriate parameter values.
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Lastly, since achieving the maximum mutation score is inherently difficult, we introduce an iterative test-generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [78, 31, 64], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [40], it handles cross-module dependencies explicitly by simply providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage; instead, it focuses on improving mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [68, 17, 65], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
To understand the contribution of each component in MutGen, we conduct an ablation study using three variants: MutGen-S (without summarization), MutGen-F (without the fixing step), and MutGen-MF (without mutation feedback). The results confirm the importance of each component in improving the overall mutation score.
In summary, this paper makes the following contributions:
  • We propose a mutation-guided, LLM-based approach (MutGen) for automatically generating test cases with high fault detection capability (i.e., mutation score), including an iterative generation process guided by mutation feedback to push the limits of LLMs in killing mutants.
  • We evaluate our approach on a widely used benchmark dataset and a new dataset we created to increase diversity.
  • We analyze the effectiveness of MutGen across different mutation operators and investigate the reasons behind the presence of live and uncovered mutants when MutGen reaches its maximum mutation score.
II Methodology
This section presents our approach MutGen through a motivating example and a formal description. We first illustrate the limitations of standard LLM-based test generation using a running example. We then precisely describe and formalize our methodology, including the complete set of user prompts employed in our approach. To save space and focus on the essential context, we omit the system prompts and the standard formatting rules widely used in existing open-source tools [48] and test generation research [24, 14, 31], which are also available in our open-source implementation.
II-A Motivating Example and Intuition
This subsection goes through a simple example to illustrate the limitations of standard prompting and to motivate our approach. The example Java code under test shown in Figure 1, which validates a date format, includes descriptive comments and input-output examples. For clarity and to save space, we retain only the most relevant parts of the comments in the figure. Most LLM-based techniques prompt LLMs using the following standard prompt:
“Generate JUnit tests for the following method to achieve high branch and line coverage.”
1package original;
2class ValidDate {
3 /** You have to write a function ......
4 * The date is valid if all of the following rules are satisfied:
5 * 1-3. ......
6 * 4. The date should be in the format: $m$-$d$-yyyy
7 * for example:
8 * validDate("03-11-2000") => True
9 * validDate("06/04/2020") => False
10 * MORE EXAMPLES ...... **/
11 public static Boolean validDate(String date) {
12 if (date.length() != 10) {
13 return false;
14 }
15 String[] dateArr = date.split("-");
16 if (dateArr.length != 3) return false;
17 int month = Integer.parseInt(dateArr[0]);
18 int day = Integer.parseInt(dateArr[1]);
19 int year = Integer.parseInt(dateArr[2]);
20 if (month < 1 $|$ month > 12) { return false; }
21 if (month == 2) {
22 if (day < 1 $|$ day > 29) { return false; }
23 } else if (month == 4 $|$ month == 6 $|$ month == 9 $|$ month == 11) {
24 if (day < 1 $|$ day > 30) { return false; }
25 } else {
26 if (day < 1 $|$ day > 31) { return false; }
27 }
28 return true;
29 }
30}
Figure 1: Example code under test
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [24, 12, 3], high coverage does not necessarily imply strong fault-detection capability when measured by the mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 1 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases, such as February 29th in leap years, likely because such examples are absent from the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats, such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the program’s comments.
Furthermore, when relying on mutation rather than code coverage, the generated inputs often fail to effectively kill mutants. For example, when the ConditionalsBoundary mutation operator in PITest replaces < with <=, it changes the conditional boundary and modifies line 24 as follows:
if (day <= 1 $|$ day > 30) return false;
This change introduces a fault: the original code accepts day == 1 as valid, while the mutant incorrectly rejects it. We observe that LLM-generated test cases, when prompted with the original code, tend to focus on representative invalid inputs (e.g., day = 0) and on valid values far from the boundary (e.g., day = 2 or higher). For instance, the model commonly produces a test such as:
assertFalse(validDate("04-00-2025"));
which fails in both the original function and the mutant. However, it frequently fails to generate:
assertTrue(validDate("04-01-2025"));
which is essential to kill the boundary-mutated version of the code.
Listing 1: Example test function generated by Llama-3.3
@Test
public void testDayBoundaryInApril() {
String date = "04-01-2025";
assertTrue(ValidDate.validDate(date));
In summary, test cases generated using the standard prompt exhibit the following limitations:
  1. 1.
    Comments in the target code can mislead LLMs, resulting in the generation of ineffective or irrelevant test cases.
  2. 2.
    Given the presence of code mutants, LLMs struggle to understand how to generate tests that kill them effectively.
To address these issues, we propose MutGen, an approach that improves the fault detection capability of LLM-generated test cases via a two-stage process: preprocessing and generation, which are detailed in Section II-C and Sections II-D-II-F, respectively.
II-B Formal Overview of MutGen
Figure 2: MutGen Overview
Let $\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$ denote the set of mutants generated from the program under test. The mutation score $\mu(T)$ of a test suite $T$ is defined as:
$\mu(\mathcal{T})=\frac{|\{m\in\mathcal{M}\mid\mathcal{T}\text{ kills }m\}|}{|\mathcal{M}|}$
We aim to generate a test suite $\mathcal{T}$ such that $\mu(\mathcal{T})$ is maximized. To achieve this, we propose a two-stage process based on prompting an LLM, namely preprocessing and generation.
The preprocessing stage prepares information to guide test case generation, including code summarization and mutation feedback, as detailed in Section II-C. In the generation stage, the LLM is prompted with mutation feedback, such as descriptions of mutation operators and diff-like representations of mutants, to generate test cases.
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{k}\}$ denote the set of mutation feedback entries (i.e., killed, live, uncovered), where each $r_{i}$ is a tuple $\langle\text{id}_{i},m_{i},s_{i},\text{op}_{i}\rangle$, respectively representing the mutant identifier, the mutated statement, the mutant status, and the applied mutation operator. MutGen first generates an initial test suite $\mathcal{T}$:
$\mathcal{T}=\text{LLM}(\text{Prompt}_{\text{mut}}(\mathcal{R}))$
Here, as further described in the following sections, $\text{Prompt}_{\text{mut}}$ is a structured prompt that encodes details of the mutation operators and specific mutant instances, as illustrated in the example $\text{Prompt}_{\text{mut}}$ shown in Figure 5.
After executing $\mathcal{T}$, MutGen collects a set $\mathcal{F}=\{f_{1},...,f_{q}\}$ consisting of all test cases that encountered execution failures and corresponding error messages. For each $f_{i}$, MutGen constructs a second prompt $\text{Prompt}_{\text{fail}}$ (as illustrated in the example shown in Figure 6) and invokes the LLM to generate a set of fixed test cases $\mathcal{T}_{fix}$:
$\mathcal{T}_{fix}=\{\text{LLM}(\text{Prompt}_{\text{fail}}(f_{i}))\mid f_{i}\in\mathcal{F}\}$
The final test suite comprises all passing test cases obtained by combining the filtered and fixed tests. We denote this using the function ExeFilter, as follows:
$\mathcal{T}_{\text{final}}=\texttt{ExeFilter}((\mathcal{T}\setminus\mathcal{F})\cup\mathcal{T}_{\text{fix}})$
This process will be detailed in Section II-E.
Our objective is to maximize the mutation score:
$\mathcal{T}^{*}=\arg\max_{\mathcal{T}}\mu(\mathcal{T})$
Our assumption, as reported in the mutation testing literature [12, 5], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [24] and therefore enables the generation of more effective test suites. However, as revealed in recent work [29], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section II-F, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
Figure 2 shows an overview of MutGen. MutGen takes a program under test as input and returns test cases aimed at improving its fault-detection effectiveness by maximizing the test suite’s mutation score. We highlight all data flows in red in the figure.
II-C Preprocessing Stage
The preprocessing stage prepares information to guide the test case generation.
Code Summarization. To address the first limitation above, MutGen removes comments from the target code and uses a prompt (shown in Figure 3) to instruct the LLM to summarize the method’s purpose and input format. The prompt used is as follows:
Prompt Used for Summarizing Summarize the following Java source code in 1-3 sentences, including its purpose and input format (if applicable). {{ source code }}
Figure 3: Prompt used for summarizing
For the target code shown in Figure 1, Llama-3.3 returns the following summary:
“This Java function, ‘validDate’, validates a given date string in the format “mm-dd-yyyy” and returns ‘true’ if the date is valid according to specific rules regarding month and day ranges, and ‘false’ otherwise. The input date string must be non-empty and follow the exact specified format. If any of these conditions are not met, the function will return ‘false’.”
Since our work focuses on improving the fault-detection capability of generated tests through mutation feedback, we directly use the summaries produced by the LLM without further optimization. Investigating how the size and quality of generated summaries affect test generation effectiveness is a promising direction for future work.
Mutation Feedback. To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with greater fault-detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [60] to generate a mutation report, as shown in Figure 4. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 5) that guides the LLM in generating more effective test cases during the generation stage.
For example, the mutation feedback extracted from the report for the mutant shown in Listing II-A is as follows:
21. conditional boundary at line 24, uncovered:
if (day<=1||day>30) return false;
This indicates that the existing tests do not cover the 21st mutant, generated by applying a conditional boundary change at line 24, specifically by replacing < with <=, as shown in the second line. Formally, this mutation feedback entry can be represented as:
$r_{21}=\langle 21,$
   $if\ (day\leq 1||day>30)\ returnfalse;$,
   uncovered,
   change conditional boundary at line 24 $\rangle$
Mutations
12 1. negated conditional -> KILLED
13 1. replaced Boolean return with True for original/ValidDate::validDate -> KILLED
16 1. negated conditional -> KILLED
2. replaced Boolean return with True for original/ValidDate::validDate -> KILLED
20 1. negated conditional -> SURVIVED Covering tests
2. changed conditional boundary -> SURVIVED Covering tests
3. changed conditional boundary-> SURVIVED Covering tests
4. replaced Boolean return with True for original/ValidDate::validDate -> SURVIVED Covering tests
5. negated conditional-> SURVIVED Covering tests
21 1. negated conditional -> SURVIVED Covering tests
22 1. negated conditional -> SURVIVED Covering tests
2. changed conditional boundary -> SURVIVED Covering tests
3. changed conditional -> SURVIVED Covering tests
4. replaced Boolean return with True for original/ValidDate::validDate -> SURVIVED Covering tests
5. changed conditional boundary-> SURVIVED Covering tests
23 1. negated conditional -> SURVIVED Covering tests
2. negated conditional -> SURVIVED Covering tests
3. negated conditional -> SURVIVED Covering tests
4. negated conditional -> SURVIVED Covering tests
24 1. negated conditional -> NO_COVERAGE
2. changed conditional boundary -> NO_COVERAGE
3. changed conditional boundary -> NO_COVERAGE
4. replaced Boolean return with True for original/ValidDate::validDate -> NO_COVERAGE
5. negated conditional -> NO_COVERAGE
26 1. changed conditional boundary -> NO_COVERAGE
2. negated conditional -> NO_COVERAGE
3. changed conditional boundary -> NO_COVERAGE
4. replaced Boolean return with True for original/ValidDate::validDate -> NO_COVERAGE
5. negated conditional -> NO_COVERAGE
28 1. replaced Boolean return with False for original/ValidDate::validDate -> NO_COVERAGE
Figure 4: Example mutation report
II-D Prompting with Mutation Feedback
To maximize the mutation score, MutGen augments the prompt with mutation feedback collected during the preprocessing stage and uses it to guide Llama-3.3 in generating test cases. Figure 5 illustrates the resulting prompt for the HumanEval-Java subject shown in Figure 1. With a single invocation of MutGen, the generated test suite contains 13 test cases and achieves a mutation score of 70% over 30 mutants.
We define the prompt construction function as follows:
$\text{Prompt}_{\text{mut}}$ $:\mathcal{R}\rightarrow\text{String},$
$\text{Prompt}_{\text{mut}}(\mathcal{R})$ $=\texttt{Format}(desc,src,\bigcup_{r_{i}\in\mathcal{R}}r_{i}).$
The function $\texttt{Format}(\cdot)$ organizes the given parameters into a structured prompt suitable for the LLM.
As illustrated in Figure 2 and the example in Section II-A, the prompt constructed by MutGen includes three components: a natural language code summarization (descdesc), the source code excluding comments (srcsrc), and the associated mutation feedback ($\mathcal{R}$). In Figure 5, the former two correspond to the prompt variables: descdesc is represented as {{summary}}, and srcsrc as {{source code without comments}}.
Regarding the content under ## Mutation Feedback, we first provide a sentence-level instruction to the LLM. The LLM is provided with live mutants and is prompted to generate tests that are more diverse for each mutant. Following this sentence-level prompt, the set of mutation feedback entries, $\mathcal{R}$, is provided as defined in Section II-C.
In addition, MutGen explicitly specifies the expected output format from the LLM, as illustrated in the prompt example shown in Figure 5.
Prompt Used for Generation # Unit Test Generator for Java using Junit4 ## Source Code Summary {{summary}} ## Source File {{source code without comments}} ## Mutation Feedback Analyze the live mutant to identify the defects; the generated test cases must look very different from the existing ones. 21. conditional boundary at line 24, uncovered: if (day<=1||day>30) return false; ## Output Format in YAML new_tests: - test_behavior: … test_name: … test_code: … new_imports: …
Figure 5: Prompt used for generation
II-E Fixing Execution Errors
After the initial generation stage, some test cases may fail due to execution errors. MutGen introduces a fixing step that leverages a structured prompt, as illustrated in Figure 6, to guide the LLM in repairing these failures. In the example shown, two generated test cases initially fail to execute due to runtime errors. By applying the fixing step, MutGen successfully repairs one of the failing test cases, resulting in a 13% increase in mutation score.
We categorize execution failures into two types:
  • Assertion Failures: Test cases that compile successfully but fail at runtime due to violated assertions or incorrect outputs.
  • Compilation Errors: Test cases that fail to compile due to syntax errors, missing types, unresolved references, or similar issues.
To address these failures, MutGen constructs a structured prompt for each failing test case:
$\text{Prompt}_{\text{fail}}(f_{i})=\texttt{Format}(t_{i},e_{i}),$
where $t_{i}$ is the original test case and $e_{i}$ is the associated compiler or runtime error message. The prompt explicitly instructs the LLM:
”You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages.”
The prompt then enumerates six targeted fix strategies:
  1. 1.
    Direct invocation errors: to address incorrect invocation of assertion functions.
  2. 2.
    assertEquals mismatches: to address oracle value inconsistencies in assertEquals assertions.
  3. 3.
    assertTrue/assertFalse errors: to address binary oracle inconsistencies in assertTrue or assertFalse assertions.
  4. 4.
    Ambiguous assertion references: to address ambiguity in assertEquals calls; explicit type casts are applied to clarify the parameter types.
  5. 5.
    Method name conflicts: to address duplicate method names within the test class.
  6. 6.
    Other runtime or compilation errors: to address errors not covered by the above strategies.
Detailed instructions for these strategies, along with a fixing example, are provided in Figure 6.
We define specialized fix functions for each failure type:
$\text{Fix}_{\text{comp}}(f_{i})$ $=\text{LLM}(\text{Prompt}_{\text{fail}}(f_{i}))$ $\text{for }f_{i}\in\mathcal{F}_{\text{comp}},$
$\text{Fix}_{\text{assert}}(f_{i})$ $=\text{LLM}(\text{Prompt}_{\text{fail}}(f_{i}))$ $\text{for }f_{i}\in\mathcal{F}_{\text{assert}}.$
The outputs of these functions are the repaired test sets:
$\mathcal{T}^{\prime}$ $=\{\text{Fix}_{\text{comp}}(f_{i})\mid f_{i}\in\mathcal{F}_{\text{comp}}\},$
$\mathcal{T}^{\prime\prime}$ $=\{\text{Fix}_{\text{assert}}(f_{i})\mid f_{i}\in\mathcal{F}_{\text{assert}}\},$
and the final set of fixed test cases is:
$\mathcal{T}_{\text{fix}}=\mathcal{T}^{\prime}\cup\mathcal{T}^{\prime\prime}.$
This prompt-based fixing strategy ensures that the LLM systematically addresses different types of execution failures, improving test execution success and contributing to higher mutation scores.
Prompt Used for Fixing ## Failed Test The following test failed in the test suite: {{ failed_test with error message }} You are processing execution failures. Please match only one of these guidelines and then try to correct the failed tests according to the given error messages. 1. Direct Invocation: • Always use assertion functions directly (e.g., assertEquals(value1, value2)) instead of prefixed calls like Assert.assertEquals(value1, value2). 2. Handling assertEquals Mismatches: • If an error follows the pattern: • test_file_name:line_number: expected:error_value but was:correct_value • And involves assertEquals, replace the expected value with the correct one. • Example: assertEquals(error_value, xxx); → assertEquals(correct_value, xxx); 3. Fixing assertTrue / assertFalse Errors: • If an error follows the pattern: • classname:line_number:null • And involves assertTrue or assertFalse, swap the function name while keeping all parameters unchanged. • Example: assertTrue(xxx); → assertFalse(xxx); 4. Handling ambiguous • If an error has the pattern: • reference to assertEquals is ambiguous • And involves assertEquals, add type casts to both parameters based on their respective types. • Example: assertEquals(expectedValue, actualValue); → assertEquals((long) expectedValue, (long) actualValue); 5. Method name conflicts • If an error follows the pattern: method is already defined in the class, • only append a number to the method name. • Example: public void test_null() → public void test_null1(); 6. Other errors • Try to fix the errors according to the error messages. ## Output Format in YAML The same applied as in Figure 5.
Figure 6: Prompt used for fixing
II-F Pushing the Limit: Iterative Generation
We propose an iterative generation process that leverages mutation feedback and failed test cases to progressively enhance the quality and effectiveness of the generated tests. At each iteration $i$, MutGen obtains a working test suite $\mathcal{T}_{i}$. Before proceeding to the next iteration, MutGen evaluates $\mathcal{T}_{i}$ to obtain mutation feedback $\mathcal{R}_{i}$. The LLM is re-invoked with $\text{Prompt}_{mut}(\mathcal{R}_{i})$ to generate an initial set of new test cases. The failing cases are filtered as $\mathcal{F}_{i}$ and, for each $f_{j}\in\mathcal{F}_{i}$, the LLM is subsequently invoked with $\text{Prompt}_{fail}(f_{j})$ to fix them. MutGen obtains $\Delta\mathcal{T}_{i}$ as the set of test cases that successfully execute. The updated test suite $\mathcal{T}_{i+1}$ is defined as:
$\mathcal{T}_{i+1}=\mathcal{T}_{i}\cup\Delta\mathcal{T}_{i},$
The process continues until convergence (i.e., no significant improvement in mutation score) or until a predefined iteration limit is reached. Once convergence is reached, the remaining live and uncovered mutants provide insights into mutation operators and program constructs that remain challenging for the LLM, which we analyze in Section III-B.
To illustrate the effectiveness of this iterative process, we revisit the running example introduced earlier. After two iterations, for the running example, MutGen achieves a 100% mutation score. In contrast, as demonstrated in Section III, for this subject in HumanEval-Java, using the vanilla prompt introduced at the beginning of this section, the LLM produces a test suite with a 53% mutation score, which remains unchanged after four iterations.
III Evaluation
As previously discussed, MutGen is designed to enhance the fault-detection capability of LLMs in automated test generation by incorporating mutation feedback. To evaluate its effectiveness, we compare MutGen with three baselines: EvoSuite, the state-of-the-art search-based technique; EvoSuitemut, a variant that replaces coverage-based fitness functions with strong mutation as described in [27]; and Genvanilla, a prompting-based approach that utilizes an LLM without incorporating mutation feedback and leveraging the comments provided in the program under test through summarization (as illustrated in Figure 5). We adopt Llama-3.3 70B as the base model in our evaluation due to its recent release, strong effectiveness on code-related tasks [30], and open-source accessibility, which ensures reproducibility and ease of deployment in both research and practical scenarios.
In addition to effectiveness comparisons, we analyze the limitations of LLM-based test generation by examining killing ratios across different mutation operators. This allows us to identify which mutant types are particularly challenging for LLMs to kill and to offer insights into potential areas for improvement.
Finally, we conduct an ablation study to evaluate the contribution of each component within MutGen to the overall effectiveness.
In sum, our evaluation addresses the following research questions.
  • RQ1: How does MutGen perform in generating test cases with high fault detection capability (i.e., mutation score) compared to EvoSuite, EvoSuitemut, and Genvanilla?
  • RQ2: Which mutation operators remain challenging for LLM-based test generation and for what reasons?
  • RQ3: What is the individual contribution of each component of MutGen to the overall effectiveness?
III-A Experiment Setup
Subjects. In our study, we primarily focus on method-level subjects, where each subject is a single Java class containing only one target method. We generate unit test cases specifically for this method. We use a widely adopted benchmark dataset that has been used in prior publications [65, 81, 68, 37], and we also create an additional dataset to enable a more diverse evaluation. More specifically, we include the following subjects:
HumanEval-Java
Following prior work [37, 65, 68, 81], we adopt the HumanEval dataset [13, 8] as one of our evaluation benchmarks. Specifically, we utilize all 160 Java subjects from the dataset. For each subject, we apply EvoSuite with its default configuration to automatically generate a test suite. We then evaluate the mutation score of the generated test suite using PITest [60]. Subjects where MutGen and EvoSuite both achieved a 100% mutation score are discarded, resulting in 104 Java subjects for our evaluation. Indeed, these subjects do not help us distinguish the two approaches.
Leetcode-Java
We search GitHub using the keyword "Leetcode Java" and select the top two repositories based on the Best match sorting criterion. These repositories contain Java solutions to algorithmic problems from LeetCode [43], which are categorized into three difficulty levels: Easy, Medium, and Hard. We exclude all solutions corresponding to Easy problems. We collect all Java source files for the remaining Medium and Hard problems and organize them into a unified package structure following the convention used in HumanEval-Java. We use the Maven build system to compile the project and remove any subjects that failed to compile. For each valid subject, we apply EvoSuite with its default settings to generate a test suite. Subjects for which the generated test suite achieved a 100% mutation score by MutGen and EvoSuite are once again discarded. Finally, for computational feasibility, we randomly select 50 subjects from each of the Medium and Hard sets, resulting in a total of 100 Leetcode-Java subjects used in our evaluation.
The average number of lines of code in HumanEval-Java methods is 41, with an average cyclomatic complexity (CC) of 4.90, which is close to the average CC of 5.46 reported for typical Java methods [52]. This indicates that HumanEval-Java exhibits representative control-flow complexity for standard Java methods. In contrast, methods in Leetcode-Java are generally more complex, with an average CC of 7.88. These statistics demonstrate that Leetcode-Java contains methods with complexity levels comparable to or exceeding those found in real-world Java code [66], making it both representative and more challenging than HumanEval-Java.
Metrics. Since our work focuses on enhancing the effectiveness of LLM-generated test cases, we adopt the mutation score, calculated as the ratio of killed mutants to the total number of mutants, as the primary evaluation metric (see Equation 1).
$\text{Mutation Score}=\frac{\text{\# killed mutants}}{\text{\# total mutants}}\times 100\%$ (1)
A mutant is considered killed if at least one generated test case causes a behavioral difference between the original program and its mutated version, typically resulting in a failed assertion or exception. In addition to the mutation score, we also report branch coverage and line coverage, in line with previous work [2, 14].
Process. To address RQ1, we use all subjects from both datasets. We begin by generating test cases with both Genvanilla and MutGen through a single end-to-end execution, recording the execution time and the number of tokens consumed. We then apply EvoSuite and EvoSuitemut to generate test cases for each subject, setting their time budget to the larger average processing times of MutGen on the two datasets. It is important to note that, unlike prior work [24], our approach generates tests entirely from scratch, without relying on any existing test cases. We evaluate each test suite using two structural coverage metrics, branch coverage and line coverage, measured by the JaCoCo tool [38]. Additionally, we use PITest [60] with its DEFAULTS mutation operator group to measure the mutation score of each test suite. The DEFAULTS group provides a set of operators commonly used in the literature [27, 71] and, according to the PITest official documentation [60], are not easy to detect and minimize the number of equivalent mutants. While PITest also offers a STRONGER operator group, it differs from DEFAULTS by only two additional mutation operators, whose applicability is rather limited. For example, the Remove Conditionals operator only targets equality checks. For the ALL group provided by PITest, the additional mutation operators generally generate mutants that are relatively easy to kill, making them less informative for our evaluation. Note that although PITest is designed for Java, its mutation operators can be applied to other languages as well [60, 71, 39]. To assess the statistical significance of the improvements achieved by our approach in terms of effectiveness, we performed non-parametric hypothesis testing using Vargha and Delaney’s A12 statistic [7, 69]. The A12 effect size was computed from the mutation score, which ranges from 0 to 1: a value of 0.5 indicates no difference (i.e., results are due to chance), whereas a value greater than 0.5 indicates that MutGen is more likely to outperform the compared approach. Conversely, a value less than 0.5 indicates the opposite. We also conducted a paired-sample Wilcoxon signed-rank test on both coverage and mutation score to evaluate whether MutGen achieves statistically significant improvements on these metrics.
To address RQ2, we analyze the mutation operators involved in the evaluation. As a preliminary experiment, we first randomly select 10 subjects from each dataset and perform seven rounds of end-to-end test generation. In each round, MutGen generates additional test cases from previously generated ones. By tracking mutation scores across iterations, we observe that they tend to converge after the fourth iteration. Therefore, we set the maximum number of repetitions to four in our experiments. We collect results at the fourth iteration, and analyze the distribution of mutation operators across all generated mutants and compute the killing ratio for each operator, that is, the proportion of killed mutants relative to the total number of mutants produced by that operator. This allows us to identify which mutation operators tend to produce mutants that are difficult or easy for LLM-generated test cases to kill. In our evaluation, mutation testing is conducted using PITest, which employs 11 mutation operators as described in [20]. For mutation operators with low killing ratios, we investigate why LLMs struggle to kill these mutants. Due to space constraints, we do not present a detailed analysis for each mutation operator or for all live and uncovered mutants. Instead, we analyze a subset of cases and identify several factors that contribute to LLMs’ inability to kill certain mutants. Nevertheless, our findings offer valuable insights and highlight important directions for future research in this area.
To address RQ3, we conduct an ablation study to evaluate the contribution of each core component in MutGen. Specifically, MutGen consists of three main components: (1) extracting and incorporating code summarization into prompts, (2) applying the fixing step to repair invalid test cases, and (3) extracting and incorporating mutation feedback into prompts. In this study, we disable one component at a time, yielding three variants: MutGen-S (using the original code comments instead of summarization), MutGen-F (omitting the fixing step), and MutGen-MF (excluding mutation feedback). For each variant, we perform four iterations of test case generation and measure mutation score to assess the impact of each component on the effectiveness of MutGen.
Implementation. We describe important implementation details of MutGen and our evaluation setup. All subjects used in the evaluation are built using the Maven build system. We follow official guidelines to integrate EvoSuite, JaCoCo, and PITest into the Maven pipeline for automated test generation, code coverage measurement, and mutation testing, respectively. Our approach, MutGen, is implemented in Python and built on top of the Mutahunter framework [48]. Specifically, we reuse components from Mutahunter for prompt handling, LLM interfacing, and response processing. We use Ollama [51] to locally deploy the Llama-3.3 70B model [30] with the default configuration, setting the temperature to 0.0 for more deterministic outputs. As our main baseline, EvoSuite officially and stably supports JUnit 4, while its support for JUnit 5 is currently limited and experimental. Therefore, JUnit 4 remains the recommended choice for reliable test generation. To ensure a fair comparison, we adopt JUnit 4 in all experimental settings.
The evaluation is conducted on a Linux server equipped with a 56-core CPU, 125 GB of RAM, a single NVIDIA RTX 6000 Ada GPU, and running Ubuntu Linux 24.04.1 LTS. Our tool and benchmarks can be found at: https://anonymous.4open.science/r/MUTGEN-4A06.
III-B Results and Analysis
TABLE I: Summary of Results: Comparison of MutGen and baseline approaches on different datasets. For each approach and dataset, we report the average per-subject line coverage (Line Cov), branch coverage (Branch Cov), mutation score (Mutation Score), number of mutants per subject (#Mutants per Subject), Vargha and Delaney’s A12 effect size (A12 effect size), total runtime in seconds (Time (secs)), and the total number of tokens consumed (#Tokens). “*” denotes cases when MutGen achieves significantly better results than the compared baselines.
Approach Dataset Line Cov Branch Cov Mutation Score #Mutants per Subject A12 effect size Time (secs) #Tokens
EvoSuite HumanEval-Java 95.6% 93.4% 69.5% 11 0.759 150 1240
Leetcode-Java 99.0% 98.9% 58.9% 19 0.899 150 956
EvoSuitemut HumanEval-Java 95.4% 93.4% 67.4% 11 0.744 150 1232
Leetcode-Java 98.1% 98.7% 58.1% 19 0.889 150 940
Genvanilla HumanEval-Java 96.2% 92.8% 77.9% 11 0.650 74.1 846
Leetcode-Java 96.3% 92.7% 69.9% 19 0.734 76.3 1232
MutGen HumanEval-Java 98.3% 95.8% 89.5%* 11 125.9 899
Leetcode-Java 98.4% 94.8% 89.1%* 19 149.4 1629
Comparison with EvoSuite, EvoSuitemut, and Genvanilla
Table I presents the overall results of EvoSuite, EvoSuitemut, Genvanilla, and our approach MutGen on the two datasets: HumanEval-Java and Leetcode-Java. The best results for each metric and dataset are highlighted.
On HumanEval-Java, for both line coverage and branch coverage, MutGen outperforms EvoSuite, EvoSuitemut and Genvanilla. On Leetcode-Java, we see the opposite: EvoSuite achieves the highest coverage scores, surpassing both LLM-based techniques. For both datasets, however, differences are relatively small, with values below 3 percentage points. Although MutGen is primarily designed to improve mutation score, the results also demonstrate that it consistently outperforms Genvanilla in terms of line and branch coverage. This supports previous observations  [24] that improvements in mutation score can also lead to enhanced code coverage.
Regarding the mutation score, PITest generates an average of 11 and 19 mutants per subject on HumanEval-Java and Leetcode-Java, respectively, resulting in a total of 1,144 and 1,900 mutants across the two datasets. MutGen achieves mutation scores of 89.5% and 89.1% on the two datasets. Compared to EvoSuite, this represents large improvements of 28.8% and 51.3%, respectively, corresponding to around 20 and 30 percentage points. Compared to Genvanilla, MutGen also significantly improves the mutation score, though to a lesser extent. Additionally, MutGen also surpasses EvoSuitemut, which specifically aims to maximize the mutation score, to an even larger extent. Indeed, surprisingly, we observe that EvoSuite slightly outperforms EvoSuitemut. This might be due to the latter being developed 10 years ago [26] and receiving no further maintenance. For example, we found that the mutation operators used by EvoSuitemut are limited, resulting in poor performance of the generated test cases when evaluated by PITest. Another tentative explanation is that, to ensure fair comparisons, recall that we set the time budget to the maximum used by MutGen, but a larger budget may be necessary for EvoSuitemut.
Furthermore, we report the A12 effect size [7] on mutation scores in the third-to-last column of Table I. For the comparison between EvoSuite and MutGen, the A12 values are 0.759 and 0.899, indicating that MutGen has a 75.9% and 89.9% probability of achieving higher mutation scores than EvoSuite on HumanEval-Java and Leetcode-Java, respectively. Similarly, compared with Genvanilla, the A12 values are 0.650 and 0.734, indicating that MutGen outperforms Genvanilla with probabilities of 65.0% and 73.4% on the same datasets. Additionally, based on the results of the paired-sample Wilcoxon signed-rank test, we found that MutGen achieves statistically significant improvements ($\alpha$ = 0.05) in mutation scores compared with EvoSuite, EvoSuitemut, and Genvanilla. However, there is no statistically significant difference in coverage, as most subjects already achieve near 100% coverage across all approaches.
In terms of overhead, MutGen consumes fewer tokens on HumanEval-Java but more tokens on Leetcode-Java. The LeetCode-Java subjects consist of solutions to medium- and hard-level algorithmic problems, with more complex logic than those in HumanEval-Java. As a result, PITest generates more mutants for them (19 vs. 11), and Llama-3.3 requires more test cases to kill all mutants. Compared with Genvanilla, EvoSuite, and EvoSuitemut, in Leetcode-Java, MutGen consumes more tokens while achieving higher mutation scores as well as better line and branch coverage. In contrast, on HumanEval-Java, MutGen consumes fewer tokens than EvoSuite and EvoSuitemut while still achieving superior performance in both coverage and mutation score. For the baseline approaches (i.e., EvoSuite and EvoSuitemut), a possible explanation for their higher token consumption on HumanEval-Java is that these programs contain more lines of code on average, whereas LeetCode-Java programs exhibit higher cyclomatic complexity. Longer programs tend to yield larger, more verbose test suites generated by these search-based techniques, resulting in increased token usage at the test level. Interestingly, we observe an opposite trend for our LLM-based approach. On LeetCode-Java, MutGen consumes more tokens than on HumanEval-Java, despite the latter having longer programs on average. We hypothesize that this behavior is related to the use of mutation feedback in the prompt. Specifically, LeetCode-Java has more mutants per subject, resulting in more uncovered or live mutants included in the mutation feedback provided to the LLM. As a consequence, the prompt becomes richer and more informative, encouraging the LLM to generate more extensive test cases to kill these mutants, which, in turn, leads to higher token consumption. However, differences of that scale in the number of tokens do not have practical implications.
Overall, the results suggest that MutGen significantly improves mutation scores relative to alternatives, with large differences, at the expense of slightly lower code coverage than EvoSuite for the most complex subject (Leetcode-Java). This is expected, as EvoSuite is designed to maximize coverage. For the other subject (HumanEval-Java), the reverse trend is observed. Even when EvoSuite is specifically configured as EvoSuitemut to maximize mutation scores [26], it performs slightly worse than EvoSuite on both datasets. Therefore, to optimize mutation scores and, thus, fault detection, we recommend using MutGen, though the impact on coverage is unclear and likely negligible in practice.
Answer to RQ1 MutGen consistently achieves higher mutation scores than EvoSuite, EvoSuitemut, and Genvanilla, demonstrating its superior fault detection capability. While this may result in a slight decrease in code coverage for more complex subjects (e.g., Leetcode-Java), the difference is practically negligible.
TABLE II: Mutant Status Counts per Operator on HumanEval-Java (Sorted by Total)
Operator K L NoCOV Total Killed Ratio
NegateConditionals 398 7 3 408 97.5%
Math 209 11 3 223 93.7%
ConditionalsBoundary 186 18 4 208 89.4%
EmptyReturns 71 1 4 76 93.4%
TrueReturns 45 7 2 54 83.3%
Increments 50 2 1 53 94.3%
PrimitiveReturns 29 4 1 34 85.3%
FalseReturns 26 0 1 27 96.3%
VoidMethodCalls 18 7 1 26 69.2%
NullReturns 19 0 0 19 100.0%
InvertNegatives 16 0 0 16 100.0%
Summary 1067 57 20 1144 93.3%
TABLE III: Mutant Status Counts per Operator on Leetcode-Java (Sorted by Total)
Operator K L NoCOV Total Killed Ratio
Math 766 8 14 788 97.2%
NegateConditionals 474 5 14 493 96.1%
ConditionalsBoundary 288 11 1 301 95.7%
PrimitiveReturns 97 14 2 113 85.8%
VoidMethodCalls 58 2 7 67 86.6%
Increments 46 7 0 53 86.8%
EmptyReturns 24 0 2 26 92.3%
TrueReturns 15 0 11 26 57.7%
NullReturns 17 0 2 19 89.5%
FalseReturns 13 0 0 13 100.0%
InvertNegatives 1 0 1 2 50.0%
Summary 1799 47 54 1900 94.7%
Effectiveness of MutGen on Different Mutation Operators
To evaluate the effectiveness and limitations of MutGen in killing mutants, we analyze the killed ratio across different mutation operators (described in [20]) after running MutGen for four iterations. Tables II and III present the mutant status counts for each mutation operator on HumanEval-Java and Leetcode-Java, respectively. In both tables, Column K indicates the number of killed mutants for each corresponding mutation operator. Column L shows the number of live mutants (covered by not killed), while Column NoCOV represents the number of mutants that are not covered by any test cases. Column Killed Ratio denotes the proportion of killed mutants relative to the total number of mutants.
We observe that, for frequently applied mutation operators (i.e., those appearing higher up in the tables), such as Math and NegateConditionals, MutGen progressively kills the corresponding mutants across iterations. After four iterations, MutGen achieves a high killed ratio—for instance, 97.2% for Math on Leetcode-Java. The remaining 2.8% of mutants are mostly uncovered, meaning they are not exercised by any generated test cases.
In contrast, several other mutation operators, including VoidMethodCalls and TrueReturns, exhibit lower effectiveness. For example, VoidMethodCalls shows a killed ratio of 69.2% on HumanEval-Java, and TrueReturns achieves only 57.7% on LeetCode-Java. Notably, although these operators differ in their occurrence frequency, our correlation analysis indicates no strong and consistent monotonic relationship between operator frequency and killing ratio. The Spearman correlation is 0.214-0.214 on HumanEval-Java and $0.36$ on LeetCode-Java. These results suggest that frequency alone does not explain the observed differences in effectiveness and that other factors, such as oracle difficulty or semantic complexity, may play a more important role.
For the uncovered mutants, the primary reasons are the presence of dead code or the inability of the generated test cases to reach the mutated statements, even after incorporating mutation feedback. This suggests that further improvements in mutation score may require more fine-grained information and targeted guidance. Exploring how to provide such detailed feedback to LLMs effectively remains an open research direction.
For live mutants, most mutation operators alter the original program logic. For example, the MATH operator may mutate a statement a = b + c; to a = b - c;, changing the value of the variable a. If a is not critical to the program’s behavior, the LLM may fail to recognize the impact of the mutation and thus generate test cases that do not expose the fault. This observation further underscores the need for finer-grained semantic information to guide LLMs in identifying and targeting meaningful behavioral changes, thereby improving their ability to detect live mutants.
Answer to RQ2 MutGen performs well on frequently applied mutation operators, achieving high killed ratios (e.g., 97.5% for NegateConditionals on HumanEval-Java), while operators that appear less frequently—such as VoidMethodCalls and TrueReturns—remain challenging, primarily due to dead code, uncovered statements, or subtle logic changes that the LLM fails to recognize, suggesting the need for more fine-grained semantic guidance.
(a) Results on HumanEval-Java
(b) Results on Leetcode-Java
Figure 7: Mutation Score Changes over 4 Iterations: Ablation Study on Both Datasets.
Ablation Study
Figure 7 presents the mutation scores achieved by the ablation variants—MutGen-S, MutGen-F, MutGen-MF—and our approach MutGen across four iterations. Specifically, Figure 7(a) shows the results on HumanEval-Java, while Figure 7(b) reports the results on Leetcode-Java. The horizontal axis represents the number of iterations, and the vertical axis denotes the mutation score.
Based on the mutation scores obtained at the fourth iteration, code summarization appears to have the most significant impact on the effectiveness of MutGen. As discussed in Section II-A, code comments can sometimes mislead LLMs or omit essential details—such as input formats—that are crucial for guiding test case generation.
The fixing stage has a notable impact on effectiveness, particularly on the LeetCode-Java dataset, where many generated test cases initially fail to execute. Prior work reports that LLM-generated tests often suffer from low execution success rates. Our results provide quantitative evidence of improvement from fixing, based on the failed test cases generated across all iterations of MUTGEN. On HumanEval-Java, MUTGEN produces 1,254 failed test cases, of which 52.08% are successfully repaired, leaving 601 unfixed test cases. On LeetCode-Java, 47.29% of the 2,742 failed test cases are fixed, leaving 1,445 unfixed tests. These repaired tests substantially contribute to improving mutation scores. We further manually inspected 50 unfixed test cases randomly sampled from the remaining failures. We found that 74% are caused by incorrect oracles (e.g., wrong assertion functions or erroneous expected outputs), while 26% are due to runtime exceptions triggered by invalid inputs (e.g., null pointer parameters). This suggests that most remaining failures stem from oracle-level semantic issues, which warrants further work on more advanced fixing strategies.
Lastly, mutation feedback also enhances effectiveness. As revealed in RQ2, although many mutants generated by common mutation operators are relatively easy for LLMs to kill, mutation feedback still yields significant gains in effectiveness by guiding the model. Future work may further investigate how mutation feedback can be tailored to help LLMs kill harder-to-detect mutants and which mutation operators benefit most from such guidance. We leave this exploration as a direction for future research.
Answer to RQ3 Among MUTGEN’s components, code summarization has the greatest impact on effectiveness by providing clear guidance to the LLM; the fixing stage increases the number of test cases that execute successfully and thereby improves the mutation score; and mutation feedback further improves performance by guiding the LLM to target harder-to-kill mutants.
III-C Threats to Validity
The primary threat to internal validity lies in the correctness of the implementation of MutGen and the experimental scripts. To reduce this threat, we have carefully reviewed our code. Another potential threat to internal validity concerns the configuration of the employed LLM, as variations in parameters such as temperature, top-p, and maximum generation length may lead to different outcomes. In our experiments, as described in Section IV.A (Implementation), we deploy Llama-3.3 70B locally using Ollama with its default configuration and set the temperature to 0.0 to obtain more deterministic outputs, keeping all other parameters fixed across all experimental settings to mitigate this threat.
Randomness may also affect the effectiveness of MutGen. However, using three repetitions in our experiments is sufficient to account for the stochasticity of LLM outputs. Empirical observations from RQ1 support this: on HumanEval-Java, MutGen achieves average mutation scores of 89.9%, 90.2%, and 88.5% across three runs (a range of 1.7%), while on Leetcode-Java, the scores are 88.2%, 88.0%, and 91.1% (a range of 2.1%). Such differences do not affect our conclusions. For the baseline generation tool, EvoSuite, we observed very similar mutation scores across three runs and all subjects in both datasets, with almost no variation. Since method-level subjects have a limited search space (around 40 lines of code per subject, as reported in Section IV.A), the evolutionary search guided by coverage converges quickly toward similar solutions across runs. Observations for EvoSuitemut, a variant of EvoSuite that maximizes mutation score, are consistent with those of EvoSuite, for the same reasons. Overall, these small differences indicate stable results across runs. Therefore, performing three repetitions is sufficient to obtain reproducible results. Reporting the average over three runs is also consistent with prior work on evaluating LLM-based approaches [78, 22, 28], including test generation. This description has been clarified in Section IV.C.
The threat to external validity mainly lies in the subjects and the target approaches. For subject selection, we reused benchmarks from existing publications and further enhanced subject diversity by evaluating our approach on 100 additional subjects. These were randomly sampled from a corpus of LeetCode algorithm solutions. In the future, we will evaluate MutGen on more complicated subjects, e.g., Defects4J. Notably, Defects4J contains real-world defects, which can provide a stronger assessment of the practical effectiveness of our approach than synthetic defects simulated via mutation testing in this study.
Regarding the baseline approaches, we selected two representative methods: (1) EvoSuite, a widely-used state-of-the-art search-based testing tool, and (2) Genvanilla, a vanilla LLM-based prompt approach. These baselines allow us to evaluate whether incorporating mutation feedback in MutGen significantly improves fault detection capability, as measured by mutation score. In addition to the above compared methods, we also compared MutGen with a variant of EvoSuite. Following published guidelines [27], we enabled the fitness function for strong mutation testing in EvoSuite. We conducted experiments on the two datasets using only the strong mutation fitness function. The average mutation scores for both datasets were 67.4% and 58.1%, respectively, which are similar to the results reported in our paper using the EvoSuite default setting. Although we use Llama-3.3 as the base LLM in this study, our approach is not tied to any specific model. MutGen is model-agnostic by design, and evaluating its performance relative to other LLMs is left for future work.
Several factors may affect the construct validity of our evaluation.
Equivalent Mutants. Our approach uses the PITest mutation testing framework, which includes mechanisms to reduce the creation of equivalent mutants. While this mitigates the problem to some extent, the remaining equivalent mutants can reduce MutGen’s efficiency and prevent it from achieving a 100% mutation score.
Oracle Problem. We assume the original code is correct and only address assertion errors that cause Maven build failures. This simplification can lead to incorrect test oracles and false positives.
Data Leakage. While some public code in benchmark datasets may appear in the LLM’s training data (e.g., Llama-3.3), the specific mutants we generate are unlikely to be memorized by the LLM. Moreover, we compare MutGen with a vanilla prompting baseline, and the observed performance gains suggest that our improvements stem from prompt engineering and strategy design rather than memorization.
IV Related Work
Mutation-Guided Test Generation. Recent studies have questioned the sufficiency of coverage as an evaluation criterion. As shown in [12, 5, 36], and pointed in recent work from Meta [3, 24], high coverage does not necessarily correlate with strong defect detection capability. Mutation testing offers a more rigorous alternative by assessing how well tests detect injected faults (mutants) [12]. Although some prior work [17, 24] has attempted to improve mutation scores during test generation, it does not fully explore the underlying limitations of LLMs in detecting and killing live mutants. Our work aims to address this gap by incorporating mutation feedback into the prompt design process for LLMs, thereby enhancing the effectiveness of the generated test cases. In addition, we introduce an iterative generation process to push the limits of LLMs in killing mutants, providing insights into their current capabilities and boundaries. This exploration also highlights opportunities for future work to guide LLMs in generating test cases with stronger defect-detection capabilities.
Other techniques. Automatic unit test generation has long been a central research topic in software engineering. Traditional efforts have primarily focused on search-based and random testing techniques, with tools such as EvoSuite [25] and Randoop [56] significantly reducing the developers’ manual testing burden. While EvoSuite was originally designed to maximize code coverage, prior work [27] has investigated enhancing mutation scores by modifying its fitness functions, an objective closely related to ours. We conducted experiments along these lines and discuss the results in Section IV.B. In short, results are not significantly different from those of the default setting of EvoSuite.
With the advent of large language models (LLMs), a new paradigm has emerged in test generation. Numerous recent surveys have summarized the rapid advancements in this area [23, 72, 34]; however, due to the large volume of related work, we only highlight a selection of representative examples in this section. Early research explored fine-tuning LLMs for code-related tasks. As LLMs became more powerful and general-purpose, however, prompt-based approaches gained traction, enabling test generation without task-specific training. Several studies [65, 77, 78, 79, 45, 61] have investigated prompt engineering strategies and mechanisms to improve test case generation effectiveness using commercial LLMs, notably OpenAI’s GPT series [1, 55, 10].
Given the high cost of commercial LLMs, and the increasing availability of high-quality open-source alternatives such as the Llama and DeepSeek series [30, 46], recent research has shifted toward optimizing prompt-based effectiveness for open-source models [54, 74, 18, 75, 35, 42]. These models are becoming popular due to their accessibility and effectiveness in code-related tasks.
LLM-based test generation has been applied across a wide range of testing scenarios. Some studies focus on multilingual test generation across languages such as Python [74, 47, 76, 68], Java [79, 74, 14, 32], and others [74, 47, 80, 15, 73]. Other works target specific testing goals, such as detecting particular bugs [47, 24], API testing [11, 19, 44, 42], or generating test oracles [41, 21, 9]. LLMs have also been integrated with traditional testing techniques, including mocking [50, 82], evolutionary algorithms [67], and program analysis [31, 59, 64, 58]. Evaluation in these works typically emphasizes code coverage metrics, such as branch and line coverage.
Unlike prior work, our approach focuses on automatically improving the mutation score with LLMs, based on mutation testing feedback, since such a score better reflects the fault-detection capability of test cases than code coverage [12, 24]. As discussed in Section II-A, high code coverage does not necessarily imply strong fault detection [12]. However, as demonstrated in our experiments (Section III-B), improving mutation score often leads to increased code coverage as a by-product.
V Discussion
We briefly discuss the limitations of our current work and outline directions for future research.
As described in Section III and IV, our experiments were conducted on independent, method-level subjects. This setup demonstrated how well LLMs perform at killing mutants when not confounded by broader program structures or dependencies. However, when transitioning to more complex datasets such as Defects4J, additional challenges arise, particularly in resolving dependencies to generate executable and meaningful test suites. We leave the investigation of such datasets for future work.
In addition, while we use Llama-3.3 as the base model in this study, MutGen is inherently model-agnostic and does not rely on any specific LLM. Although our results demonstrate the effectiveness of the approach, other models may yield even better results. A more comprehensive evaluation across different LLMs is also left as future work.
Finally, although MutGen achieves high mutation scores on the selected subjects, there remain open challenges in further improving these scores. For example, guiding LLMs to effectively kill mutants produced by certain mutation operators, such as VoidMethodCalls, remains difficult. In our evaluation on HumanEval-Java, this operator achieved only a 69.2% kill ratio. Developing strategies to better handle such challenging mutants is an important direction for future research. We also observe that some generated test cases that are not repaired by the fixing stage are close to killing hard-to-detect mutants, but fail due to issues such as incorrect assertions or incomplete oracles. As analyzed in Section III.B(c), 74% of the unfixed test cases are caused by incorrect or incomplete oracles, while the remaining 26% fail due to runtime exceptions (e.g., invalid inputs). This indicates that these unfixed tests warrant more complex oracle repair strategies. In addition, an important avenue for future work is to more efficiently characterize the upper bound of LLM mutation scores across diverse target programs. In particular, identifying how quickly an LLM approaches its performance limit and how this limit varies across program sizes, mutation operators, and code structures remains an open problem.
VI Conclusion
In this paper, we propose MutGen, a mutation-guided, LLM-based approach for automatically generating unit tests with high fault detection capability. The core innovation of MutGen lies in incorporating mutation feedback into the prompt to help LLMs generate tests that kill more mutants. In addition, two components—code summarization and code fixing—further enhance the approach’s effectiveness.
We evaluated MutGen on 104 subjects from the widely-used HumanEval-Java dataset and 100 subjects from Leetcode-Java, which we constructed from LeetCode algorithmic problem solutions. Experimental results show that MutGen significantly outperforms both EvoSuite and Llama-3.3 using a vanilla prompting strategy, achieving mutation scores of 89.5% and 89.1% on HumanEval-Java and Leetcode-Java, respectively. Since MutGen aims to maximize the mutation score of generated test cases, we also compare it with a variant of EvoSuite (EvoSuitemut), where the fitness function of maximizing coverage is replaced with one that maximizes mutation score. The results show that EvoSuitemut performs similarly to EvoSuite, but both are outperformed by MutGen.
Moreover, MutGen employs an iterative generation process to push the limits of LLMs in killing mutants. We further analyze the causes of remaining live and uncovered mutants and the impact of different mutation operators, offering insights and directions for future research.